Find the points of intersection of the line y=x+1 and x^2+y^2=25 algebraically .
Answer by Boreal(15235)

y=x+1
x^2+y^2=25
substitute
x^2+(x+1)^2=25
x^2+x^2+2x+1=25
2x^2+2x-24=0
x^2+x-12=0
(x+4)(x-3)=0
x=3 or -4
y=4 or -3
(3, 4) (-3, -4)